### v1

@@ 4.4 Implementation
  • add_lemma($lemma$): Adds the given $lemma$ to the solver. GipSAT will maintain the lemmas of a frame. The add_clause method is replaced by this, as users only need to add lemmas after creating the GipSAT instance.
  • solve($assumption,constraint,droot$): Checks the satisfiability under the specified $assumption$ and $constraint$ within the domain $\text{COI}(droot)\cup\mathcal{V}(F)$. Users should ensure that $\mathcal{V}(assumption)\cup\mathcal{V}(constraint)\subseteq\text{COI}(droot)\cup\mathcal{V}(F)$.
@@
  • set_domain($droot$): Configures GipSAT to use $\mathcal{V}(F_{i})\cup\text{COI}(droot)$ as the domain for subsequent solving processes, bypassing the need to recompute the domain before each solving step.
@@
  • relind($c$): Checks the satisfiability of $Q_{relind}$.
  • has_bad(): Checks the satisfiability of $Q_{bad}$.
@@ 5.2 Experimental Results
Table 2 and Fig. 4 respectively represent the summary results of different configurations and the number of cases solved over time by different solvers in and rIC3. It can be observed that GipSAT solved $23$ more cases compared to the best-performing regular solver, Minisat, significantly enhancing the capability of the IC3 algorithm. The last column of Table 2 lists the geometric mean solving time speedup of GipSAT relative to regular solvers. It can be observed that GipSAT has achieved a significant improvement in performance, with an average speedup of $3.61$ times in solving time compared to Minisat in rIC3. When integrated with GipSAT, rIC3 solved more cases compared to the IC3 implementations in state-of-the-art model checkers. Fig. 5 presents scatter plots that compare the solving times of GipSAT with those of regular solvers in rIC3. It is quite evident that, when using GipSAT, the solving time for the vast majority of cases has significantly decreased. This experimental result clearly demonstrates the effectiveness of GipSAT in enhancing the efficiency and scalability of the IC3 algorithm.
Table 2: Total number of cases solved for different configurations. The additional cases solved and the geometric mean of the solving time ratio relative to GipSAT (the baseline) are presented, along with the PAR-2 score for each configuration.

### v2

@@ 5 Implementation
  • add_lemma($lemma$): Adds the given lemma to the solver. GipSAT maintains the lemmas of each frame. The add_clause method has been replaced with this, as users only need to add lemmas after creating a GipSAT instance. A lemma is a clause that involves variables in $X$ and has no dependent variables. Adding a clause with dependent variables could affect the correctness of domain maintenance. Therefore, only the add_lemma interface is provided.
  • solve($assumption,constraint,droot$): Checks the satisfiability under the specified $assumption$ and $constraint$ within the domain $\text{COI}(droot)\cup\mathcal{V}(F)$, where $F$ represents the lemmas maintained by GipSAT. Users should ensure that $\mathcal{V}(assumption)\cup\mathcal{V}(constraint)\subseteq\text{COI}(droot)\cup\mathcal{V}(F)$.
@@
  • set_domain($droot$): Configures GipSAT to use $\mathcal{V}(F)\cup\text{COI}(droot)$ as the domain for subsequent solving processes, bypassing the need to recompute the domain before each solving step.
@@
  • relind($c$): Checks the satisfiability of $Q_{relind}$, which is functionally equivalent to solve($\lnot c^{\prime},c,\mathcal{V}(c)\cup\mathcal{V}(c^{\prime})$).
  • has_bad(): Checks the satisfiability of $Q_{bad}$, which is functionally equivalent to solve($P,true,\mathcal{V}(P)$).
@@ 6.2 Experimental Results
Table 2 and Fig. 4 present the summary of results and the number of cases solved over time by various configurations. The comparison shows that rIC3 performs well relative to state-of-the-art systems like ABC and nuXmv, making it a suitable choice as the experimental foundation. Notably, GipSAT solved $23$ more cases than the best-performing regular solver, Minisat, greatly enhancing the capability of the IC3 algorithm. The last column of Table 2 lists the geometric mean speedup in solving time of GipSAT relative to regular solvers. GipSAT achieves a significant improvement in efficiency, with an average speedup of $3.61$ times compared to Minisat in rIC3. Fig. 5 shows scatter plots comparing the solving times of GipSAT with those of regular solvers in rIC3. It is clear that, with GipSAT, the solving time for most cases has decreased substantially. This experimental result clearly highlights the effectiveness of GipSAT in improving the efficiency and scalability of the IC3 algorithm.
Table 2: Total number of cases solved for different configurations. The additional cases solved and the geometric mean of the solving time ratio relative to GipSAT are presented, along with the PAR-2 score for each configuration.